. An object moves along a straight line with an acceleration of 2m/s2. If its initial speed is 10m/s, what will be its speed 5s later? 

Grade:9

4 Answers

VINEETH N
8 years ago
zero
 
Neeti
8 years ago
 
initial velocity u = 10m/s , acceleration = 2m/s2 , time t = 5seconds.
A formula in 1D is v = at where v is the final velocity of the body.
therefore,
= 10 + 2 x 5
   = 20m/s
 
Please approve the question if you like it. Feel free to ask any follow-up question. :)
 
Since you’ve mentioned acceleration so the equation uses +at , If it had been deccelaration then v = u- at
VINEETH N
8 years ago
I thought the formula of a=v-u/t. I think my answer was actually wrong
 
Neeti
8 years ago
Yes your formula is incorrect. it’s v = u + at for accelaration and v = u – at for deccelaration